OpenStudy (anonymous):

#2 Determine whether the graph of y = −2x2 − 3x + 10 opens up or down and whether it has a maximum or minimum point. Down and Maximum Down and Minimum Up and Maximum Up and Minimum

OpenStudy (anonymous):

This is from the family of functions, y=x^2 which are parabolas that open vertically. If the x^2 term is negative, then it's a frowny face (opens down). From that shape, what can you say about whether it has a max or min?

OpenStudy (anonymous):

up and max?

OpenStudy (anonymous):

nope, the opposite. down and max

OpenStudy (anonymous):

also this 1 please Identify which equation is a quadratic equation. y = −5x − 3 y = −4x(x + 10) + 34 y + x2 = (x − 4)(x + 3) y − 2 = (x − 2)(x2 + 1)

OpenStudy (anonymous):

Simplify them and see which one has a variable with and exponent of 2.